savannalily:

tan x° = e divided by 8 and sin x° = e divided by f

savannalily:

Look at the figure: An image of a right triangle is shown with an angle labeled x. If tan x° = e divided by 8 and sin x° = e divided by f, what is the value of cos x°? cos x° = 8 divided by f cos x° = f divided by 8 cos x° = 8f cos x° = 8e

Vocaloid:

tan(x) = opposite/adjacent = e/8 sin(x) = opposite/hypotenuse = e/f identify your opposite and hypotenuse based these equalities. Then plug them into cos(x) = adjacent/hypotenuse Alternatively, tan(x) = sin(x)/cos(x) and you could calculate cos this way